OpenStudy (anonymous):

The function f(x) = x². The graph of g(x) is f(x) translated to the right 3 units and down 3 units. What is the function rule for g(x)?

OpenStudy (alekos):

that would be g(x) = f(x-a) + b where a is the number of units shifted in the x-axis and b is the number of units shifted in the y-axis

OpenStudy (anonymous):

g(x) = f(x-3) - 3 ?

OpenStudy (alekos):

that's it, you got it!

OpenStudy (anonymous):

Thanks :) If we had g(x) = f(x+3) - 3 The plus three moves it to the left right?

OpenStudy (alekos):

f(x-3) shifts it in the positive x direction by 3 units

OpenStudy (alekos):

f(x+3) shifts it in the negative x direction by 3 units

OpenStudy (alekos):

so your right

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (alekos):

no problem